OpenStudy (v-sync):

Find the sum of the measures of the interior angles of a convex 30-gon. give answer and show all work.

OpenStudy (mathstudent55):

The sum of the measures of the interior angles of an n-sided convex polygon is \((n - 2)180\)

OpenStudy (v-sync):

OK so (n - 2)180 = (30 - 2)180 = (28)180 = 5040 correct?

OpenStudy (mathstudent55):

Correct. This formula is easy to remember.

OpenStudy (mathstudent55):

Use 2 less than the number of sides and multiply by 180. Why 180? It's because 180 is the sum of the measures of the interior angles of a triangle.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Notice that for each polygon, the number of triangles you can divide it into is always 2 fewer than the number of sides. Since each triangle has a sum of measures of angles of 180, multiply 2 fewer than the number of sides to find the sum of the measures of the angles of a polygon.

OpenStudy (mathstudent55):

It is easier to remember the formula if it makes sense to you. S = (n - 2)180

OpenStudy (v-sync):

Thanks